# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ Abstract
Since their introduction, Kolmogorov–Arnold Networks (KANs) have been successfully applied across several domains, with physics-informed machine learning (PIML) emerging as one of the areas where they have thrived. In the PIML setting, Chebyshev-based physics-informed KANs (cPIKANs) have become the standard due to their computational efficiency. However, like their multilayer perceptron-based counterparts, cPIKANs face significant challenges when scaled to depth, leading to training instabilities that limit their applicability to several PDE problems. To address this, we propose a basis-agnostic, Glorot-like initialization scheme that preserves activation variance and yields substantial improvements in stability and accuracy over the default initialization of cPIKANs. Inspired by the PirateNet architecture, we further introduce Residual-Gated Adaptive KANs (RGA KANs), designed to mitigate divergence in deep cPIKANs where initialization alone is not sufficient. Through empirical tests and information bottleneck analysis, we show that RGA KANs successfully traverse all training phases, unlike baseline cPIKANs, which stagnate in the diffusion phase in specific PDE settings. Evaluations on seven standard forward PDE benchmarks under a fixed training pipeline with adaptive components demonstrate that RGA KANs consistently outperform parameter-matched cPIKANs and PirateNets – often by several orders of magnitude – while remaining stable in settings where the others diverge.
@@ 1 Introduction
The widespread adoption and integration of machine learning into computational science has profoundly influenced the way complex physical phenomena are modeled and analyzed. One of the most striking advances is the Physics‑Informed Machine Learning (PIML) framework [1, 2], which offers a compelling alternative to traditional discretization‑based solvers for both forward and inverse problems involving partial differential equations (PDEs). Within the PIML framework, the governing equations, alongside boundary and initial conditions, plus any observational data, are embedded into a differentiable loss function, while a neural network parametrizes the unknown solution field. Leveraging automatic differentiation [3] to evaluate differential operators exactly, PIML eliminates the need for mesh generation and yields continuous, high‑fidelity predictions with reduced computational cost. As a result, it has found success across a broad spectrum of scientific and engineering disciplines, from fluid mechanics [4, 5, 6, 7] and materials science [8, 9] to medicine [10, 11, 12] and chemistry [13, 14].
@@
While a variety of neural architectures have been explored within the PIML paradigm, including convolutional neural networks (CNNs) [15], generative adversarial networks (GANs) [16], and long short-term memory (LSTM) networks [17], the fully connected multilayer perceptron (MLP) is the predominant backbone. When an MLP parametrizes the solution field, the variant is conventionally termed a Physics-Informed Neural Network (PINN), which is also the original formulation of PIML [18]. Despite their widespread use, PINNs exhibit several well‑documented shortcomings, including spectral bias toward low‑frequency modes [19], restricted interpretability and limited scalability with depth, among other challenges in their training dynamics. To address these issues, numerous mitigation strategies have been proposed, ranging from architectural modifications [20, 21, 22, 23], to adaptive training techniques [24, 25, 26]. A complementary approach is to forgo the MLP backbone altogether in favor of alternative architectures that mitigate several of these issues by design.
@@
One such emerging alternative is the Kolmogorov–Arnold Network (KAN) [27]. Whereas the expressivity of MLPs is supported by the universal approximation theorem [28], KANs are grounded in the Kolmogorov–Arnold representation theorem [29]. In practice, a KAN layer replaces fixed nonlinear activations with a learnable linear combination of basis functions; the original implementation employs B-splines, but other basis functions such as radial basis functions [30], Chebyshev polynomials [31] and Rectified Linear Unit (ReLU)-based functions [32] have also been explored. This design offers several benefits, most notably in terms of enhanced interpretability [33] and robustness against spectral bias [34]. These have motivated the development of Physics-Informed Kolmogorov–Arnold Networks (PIKANs) [35, 36], where the MLP backbone is substituted by a KAN in the PINN framework. Initial studies have demonstrated that PIKANs can attain higher accuracy on benchmark PDEs, or comparable accuracy with considerably smaller network architectures than their MLP-based counterparts [27, 35, 37]. Consequently, they have already seen practical deployment in a variety of scientific and engineering contexts [38, 39, 40].
@@
Despite their promising early results, PIKANs present their own challenges. Computational overhead is the most immediate: evaluating and differentiating the B‑spline basis in the original KAN formulation quickly becomes a bottleneck, leading most physics-informed implementations to adopt the more efficient Chebyshev variant (cPIKAN) [35]. Scalability is another concern; empirical studies report training instabilities as the number of network parameters increases beyond a point [41, 42], limiting applicability in deep learning regimes. Similar issues are observed in deep PINNs, although the PirateNet architecture appears to mitigate them [23]. Methodological gaps also persist. Weight initialization schemes are still largely ad hoc: each basis family (B‑spline, radial‑basis, Chebyshev, etc.) provides its own default, yet no analogue to the well-studied Glorot initialization for PINNs [43] has been empirically or theoretically analyzed. Systematic experimentation with initialization schemes is at a nascent stage and has so far concentrated exclusively on B-spline-based KANs [44]. The picture is similar for adaptive training strategies. While several PINN-oriented techniques have been ported to PIKANs [37, 45, 46], a unified training pipeline, comparable to the one codified for PINNs [47], has yet to be established.
@@
Building on this foundation, we address the depth-scaling issue. Because each KAN layer carries multiple learnable basis coefficients, a KAN layer of the same width as an MLP layer is substantially more parameter-heavy; real-world tasks that need larger capacity must therefore rely on greater depth, which in turn demands stable training. To this end, we introduce a Residual-Gated Adaptive KAN (RGA KAN) architecture. We analyze its training dynamics through the lens of the Information Bottleneck (IB) framework [48] and empirically demonstrate that RGA KANs remain stable and train effectively at depths where baseline cPIKANs diverge.
@@
The remainder of this paper is structured as follows. Section 2 reviews the theoretical foundations of our study, covering the PIML framework, KANs, and the adaptive training methods applied herein. Section 3 presents the proposed basis-agnostic Glorot-like initialization and demonstrates its clear advantage over the default cPIKAN initialization through small-scale function-fitting and PDE benchmarks. Section 4 addresses the depth-scaling limitations of cPIKANs by introducing the RGA KAN architecture; we analyze its training dynamics via the IB theory and show that it remains stable where cPIKANs diverge. Section 5 delivers a comprehensive empirical comparison among RGA KANs, baseline cPIKANs, and PirateNets on a suite of forward PDE problems, supplemented by ablation studies that isolate the contribution of each adaptive training component. Finally, Section 6 summarizes our principal findings and outlines promising directions for future research.
@@ 2.1 Problem Formulation
where $\lVert\cdot\rVert_{2}$ denotes the $L^{2}$ norm and $\left\{\left(t_{\xi}^{i},\mathbf{x}_{\xi}^{i}\right)\right\}_{i=1}^{N_{\xi}}$, with $\xi$ being either “pde”, “ic” or “bc”, correspond to collocation points used to calculate the PDE’s, initial condition’s and boundary conditions’ residuals, respectively. We remark that the global $\lambda$-weights can either be defined based on domain knowledge (e.g., [35]), or adjusted dynamically during the network’s training [24, 25, 42, 49]. Moreover, we note that the set of collocation points used to calculate the PDE’s residuals can be sampled once from a fixed grid or adaptively re-sampled throughout training [50, 51, 52, 53].
@@ 2.2 Kolmogorov–Arnold Networks
Inspired by the Kolmogorov–Arnold representation theorem [29], the authors of [27] introduced KANs, a novel class of neural networks that have since been adopted as an alternative to MLPs. The formulation corresponding to Eq. (10) for the original implementation of KANs, known as “vanilla” KANs, is given by
@@
is a residual function and $B_{m}^{(l)}\left(\cdot\right)$ are univariate spline basis functions. The superscript $(l)$ reflects the inherent dependency of the spline basis functions on a layer-specific grid, while the subscript $m$ runs from 1 to $D=G+k$, where $G$ is the number of grid intervals and $k$ is the order of the basis functions. Comparing Eq. (13) to Eq. (10), a fundamental distinction arises between MLPs and KANs in terms of their functional representation: in MLPs, nonlinearity is introduced through fixed activation functions, while only the linear transformations between layers are trainable; in contrast, KANs replace these static activation functions with learnable ones, an aspect that highlights their potential to create more expressive architectures [27, 34].
While vanilla KANs have demonstrated promising results in solving forward PDE problems [36, 37], their training is burdened by the expensive computation of spline basis functions. Additionally, their dependency on a grid, although beneficial in certain applications [27, 33, 37], becomes redundant when the grid remains fixed throughout training. To mitigate these inefficiencies, some approaches have introduced optimization strategies [54, 55], while others have explored alternative, more computationally efficient basis functions [30, 31, 56]. In this work, we adopt the latter approach and use Chebyshev-based KANs, due to their proven success in PIML [35, 40, 42, 57]. We therefore modify the expression of Eq. (13) to
@@
where $T_{m}\left(\cdot\right)$ are Chebyshev polynomials of the first kind. The hyperbolic tangent in the argument of $T_{m}$ maps $x$ to the [1,1][-1,1] range, where the absolute value of the Chebyshev polynomials is bounded by 1. For the purposes of the present work, the Chebyshev polynomials are explicitly defined as functions up to order $D$ to maximize computational efficiency.
@@
2.3 Adaptive Training Methods
2.3.1 Collocation Points Resampling
The selection of collocation points used to enforce the PDE can be performed in a one-time manner, where a fixed set of training points is used throughout the entire training process, either with or without mini-batching. However, periodically resampling the collocation points during training can serve as a regularization technique, improving the network’s ability to generalize to spatiotemporal regions that were not explicitly sampled. Moreover, if the resampling process is not purely random but instead adaptive – e.g., guided by residuals or other heuristics – it has been shown to significantly enhance the final accuracy of the trained network [50, 52, 53].
For this study, we adopt the Residual-Based Adaptive Distribution (RAD) technique [51]. In particular, an initial dense pool of $N_{\text{pool}}$ collocation points, $\left\{\left(t_{\text{pool}}^{i},\mathbf{x}_{\text{pool}}^{i}\right)\right\}_{i=1}^{N_{\text{pool}}}$, is generated from a uniform grid, and training is performed on a dynamically selected subset of $N_{\text{pde}}\ll N_{\text{pool}}$ points. These points are periodically resampled from the pool according to the probability density function
2.3.2 Global Loss Weighting
In this work, we address this issue using the learning-rate annealing algorithm introduced in [20]. During training, the weight adjustment is guided by the computation
[…]
2.3.3 Causal Training
A common challenge in training neural networks to solve time-dependent PDEs is the violation of causality. Since collocation points are sampled from the entire temporal domain, the network may unintentionally minimize residuals associated with future states before adequately minimizing those corresponding to past states. To mitigate this issue, following [25], we partition the temporal domain into $M$ sequential segments of equal length and introduce temporal weights $\left\{w_{i}\right\}_{i=1}^{M}$, modifying Eq. (7) as
2.3.4 Residual-Based Attention
The learning-rate annealing algorithm introduced in Section 2.3.2 is essentially a global loss weighting scheme, while causal training corresponds to a batch-wise weighting strategy, where groups of collocation points share the same weights. However, point-wise multipliers have also demonstrated significant success in PIML [58] and have been integral to works achieving state-of-the-art results [42, 59]. In such methods, the individual loss terms in Eqs. (7)-(9) are modified as
For this study, we adopt the Residual-Based Attention (RBA) method introduced in [26], where all local weights are initially set to 1 and updated at each training iteration according to
where $\gamma\geq 0$, $\eta\geq 0$ are hyperparameters of the method.
In theory, while RBA is not inherently incompatible with causal training – as the batch-wise weights can be computed using the modified loss of Eq. (23) – the same does not hold for RAD as presented in 2.3.1. Since RAD involves resampling collocation points, it removes the one-to-one correspondence between them and their local weights. In order to apply these methods in conjunction, we adopt the following strategy: instead of assigning local weights solely to the currently selected $N_{\text{pde}}$ training points, we define the RBA weights over the entire pool, i.e., $\alpha_{\text{pool}}^{i}=1$ for $i=1,\dots,N_{\text{pool}}$ at initialization. These weights are updated at each training iteration according to Eq. (24), using the residuals evaluated at the corresponding points. Then, during each application of RAD, the probability density function in Eq. (18) is modified by incorporating the RBA weights as multiplicative factors on the residuals:
@@ 3 KAN Initialization
Initialization plays a critical role in the training dynamics of deep neural networks. In the case of MLPs, extensive theoretical and empirical work has led to well-established initialization schemes tailored to different architectures and activation functions. Within the PINN framework, architectures are most often initialized using the Glorot scheme [47], which aims to preserve the variance of activations and gradients across layers to avoid vanishing or exploding signals, thereby enabling stable deep learning.
For KANs – and, by extension, PIKANs – the choice of weight initialization remains largely ad hoc, with each variant in the literature adopting its own heuristic procedure. In contrast to the extensive body of work on MLP initialization, systematic studies for KANs are scarce and have thus far been limited to the vanilla formulation with B-spline bases [44], where power-law and LeCun-like schemes were proposed. Building on these early efforts, we develop a Glorot-like initialization derived from a variance-preservation analysis and formulated to be agnostic to the choice of basis functions. In the present work, we assess its performance in Chebyshev-based KANs and cPIKANs, demonstrating that it offers improved training stability and accuracy over the current standard initialization scheme.
@@ 3.1 Proposed Scheme
A central principle in initialization design, which is also the core idea behind the LeCun initialization [60], is to preserve the variance of the signal during the forward pass. Applying this condition to Eq. (26) gives
@@
In his original work, Glorot [43] further required variance preservation during the backward pass, so that gradients propagate across layers without amplification or attenuation. Applying the same reasoning to the gradient signal through Eq. (26) leads to
@@
This expression defines a basis-agnostic Glorot-like initialization rule: $\mu_{m}^{(0)}$ and $\mu_{m}^{(1)}$ capture the effect of the chosen basis functions on variance, while $d_{\text{I}}$ and $d_{\text{O}}$ play the same role as in the original Glorot scheme. The additional factor $1/D$ accounts for the contribution of the $D$ basis terms associated with each input dimension. In the special case $\mu_{m}^{(0)}=\mu_{m}^{(1)}=1$ and $D=1$, corresponding to the MLP setting, which can be interpreted as using a single basis function, Eq. (32) reduces exactly to the standard Glorot initialization. Detailed derivations of Eqs. (28) and (30) are provided in A.1.
For practical applications, it is often convenient to introduce a multiplicative gain factor, following common practice in initialization utilities provided by deep learning frameworks such as PyTorch [61]. This gain term allows for empirical correction in cases where the input distribution deviates from the unit-variance assumption. Incorporating this factor, the final initialization rule becomes

### v2

# Training deep physics-informed Kolmogorov–Arnold networks

@@ Abstract
Since their introduction, Kolmogorov–Arnold Networks (KANs) have been successfully applied across several domains, with physics-informed machine learning (PIML) emerging as one of the areas where they have thrived. In the PIML setting, Chebyshev-based physics-informed KANs (cPIKANs) have become the standard due to their computational efficiency. However, like their multilayer perceptron-based counterparts, cPIKANs face significant challenges when scaled to depth, leading to training instabilities that limit their applicability to several PDE problems. To address this, we propose a basis-agnostic, Glorot-like initialization scheme that preserves activation variance and yields substantial improvements in stability and accuracy over the default initialization of cPIKANs. Inspired by the PirateNet architecture, we further introduce Residual-Gated Adaptive KANs (RGA KANs), designed to mitigate divergence in deep cPIKANs where initialization alone is not sufficient. Through empirical tests and information bottleneck analysis, we show that RGA KANs successfully traverse all training phases, unlike baseline cPIKANs, which stagnate in the diffusion phase in specific PDE settings. Evaluations on nine standard forward PDE benchmarks under a fixed training pipeline with adaptive components demonstrate that RGA KANs consistently outperform parameter-matched cPIKANs and PirateNets – often by several orders of magnitude – while remaining stable in settings where the others diverge.
@@ 1 Introduction
The widespread adoption and integration of machine learning into computational science has profoundly influenced the way complex physical phenomena are modeled and analyzed. One of the most striking advances is the Physics‑Informed Machine Learning (PIML) framework [NaturePIML, ReviewPIML], which offers a compelling alternative to traditional discretization‑based solvers for both forward and inverse problems involving partial differential equations (PDEs). Within the PIML framework, the governing equations, alongside boundary and initial conditions, plus any observational data, are embedded into a differentiable loss function, while a neural network parametrizes the unknown solution field. Leveraging automatic differentiation [AD] to evaluate differential operators exactly, PIML eliminates the need for mesh generation and yields continuous, high‑fidelity predictions with reduced computational cost. As a result, it has found success across a broad spectrum of scientific and engineering disciplines, from fluid mechanics [fluid1, fluid2, fluid3, fluid4] and materials science [mat1, mat2] to medicine [med1, med2, med3] and chemistry [chem1, chem2].
@@
While a variety of neural architectures have been explored within the PIML paradigm, including convolutional neural networks (CNNs) [archcnn], generative adversarial networks (GANs) [archgan], and long short-term memory (LSTM) networks [archlstm], the fully connected multilayer perceptron (MLP) is the predominant backbone. When an MLP parametrizes the solution field, the variant is conventionally termed a Physics-Informed Neural Network (PINN), which is also the original formulation of PIML [Raissi]. Despite their widespread use, PINNs exhibit several well‑documented shortcomings, including spectral bias toward low‑frequency modes [SpectralBias], restricted interpretability and limited scalability with depth, among other challenges in their training dynamics. To address these issues, numerous mitigation strategies have been proposed, ranging from architectural modifications [Wang1, Wang2, RWF, PirateNets], to adaptive training techniques [WangNTK, Causal, RBA]. A complementary approach is to forgo the MLP backbone altogether in favor of alternative architectures that mitigate several of these issues by design.
@@
One such emerging alternative is the Kolmogorov–Arnold Network (KAN) [KAN1]. Whereas the expressivity of MLPs is supported by the universal approximation theorem [UAT], KANs are grounded in the Kolmogorov–Arnold representation theorem [KART]. In practice, a KAN layer replaces fixed nonlinear activations with a learnable linear combination of basis functions; the original implementation employs B-splines, but other basis functions such as radial basis functions [rbfKAN], Chebyshev polynomials [chebyKAN] and Rectified Linear Unit (ReLU)-based functions [reluKAN] have also been explored. This design offers several benefits, most notably in terms of enhanced interpretability [KAN2] and robustness against spectral bias [KAN3]. These have motivated the development of Physics-Informed Kolmogorov–Arnold Networks (PIKANs) [FAIR, KINNs], where the MLP backbone is substituted by a KAN in the PINN framework. Initial studies have demonstrated that PIKANs can attain higher accuracy on benchmark PDEs, or comparable accuracy with considerably smaller network architectures than their MLP-based counterparts [KAN1, FAIR, Rigas]. Consequently, they have already seen practical deployment in a variety of scientific and engineering contexts [pikanapp1, pikanapp2, pikanapp3].
@@
Despite their promising early results, PIKANs present their own challenges. Computational overhead is the most immediate: evaluating and differentiating the B‑spline basis in the original KAN formulation quickly becomes a bottleneck, leading most physics-informed implementations to adopt the more efficient Chebyshev variant (cPIKAN) [FAIR]. Scalability is another concern; empirical studies report training instabilities as the number of network parameters increases beyond a point [KANlimit, KKANs], limiting applicability in deep learning regimes. Similar issues are observed in deep PINNs, although the PirateNet architecture appears to mitigate them [PirateNets]. Methodological gaps also persist. Weight initialization schemes are still largely ad hoc: each basis family (B‑spline, radial‑basis, Chebyshev, etc.) provides its own default, yet no analogue to the well-studied Glorot initialization for PINNs [Glorot] has been empirically or theoretically analyzed. Systematic experimentation with initialization schemes is at a nascent stage and has so far concentrated exclusively on B-spline-based KANs [MOSS]. The picture is similar for adaptive training strategies. While several PINN-oriented techniques have been ported to PIKANs [Rigas, AAH1, AAH2], a unified training pipeline, comparable to the one codified for PINNs [ExpertsGuide], has yet to be established.
@@
Building on this foundation, we address the depth-scaling issue. Because each KAN layer carries multiple learnable basis coefficients, a KAN layer of the same width as an MLP layer is substantially more parameter-heavy; real-world tasks that need larger capacity must therefore rely on greater depth, which in turn demands stable training. To this end, we introduce a Residual-Gated Adaptive KAN (RGA KAN) architecture. We analyze its training dynamics through the lens of the Information Bottleneck (IB) framework [IBTheory] and empirically demonstrate that RGA KANs remain stable and train effectively at depths where baseline cPIKANs diverge.
@@
The remainder of this paper is structured as follows. Section 2 reviews the theoretical foundations of our study, covering the PIML framework and KANs. Section 3 presents the proposed basis-agnostic Glorot-like initialization and demonstrates its clear advantage over the default cPIKAN initialization through small-scale function-fitting and PDE benchmarks. Section 4 addresses the depth-scaling limitations of cPIKANs by introducing the RGA KAN architecture; we analyze its training dynamics via the IB theory and show that it remains stable where cPIKANs diverge. Section 5 delivers a comprehensive empirical comparison among RGA KANs, baseline cPIKANs, and PirateNets on a suite of forward PDE problems, supplemented by ablation studies that isolate the contribution of each adaptive training component. Finally, Section 6 summarizes our principal findings and outlines promising directions for future research.
@@ 2.1 Problem Formulation
where $\lVert\cdot\rVert_{2}$ denotes the $L^{2}$ norm and $\left\{\left(t_{\xi}^{i},\mathbf{x}_{\xi}^{i}\right)\right\}_{i=1}^{N_{\xi}}$, with $\xi$ being either “pde”, “ic” or “bc”, correspond to collocation points used to calculate the PDE’s, initial condition’s and boundary conditions’ residuals, respectively.
While the standard formulation described above relies on fixed weights and static sets of collocation points, many practical applications require adaptive training strategies. The framework can thus be extended by introducing dynamic global or local weights, as well as adaptive resampling schemes. In this work, we specifically incorporate a suite of such adaptive strategies, namely residual-based attention (RBA), residual-based adaptive distribution (RAD), causal training and learning-rate annealing (LRA). A detailed description of these methods is provided in A.
@@ 2.2 Kolmogorov–Arnold Networks
Inspired by the Kolmogorov–Arnold representation theorem [KART], the authors of [KAN1] introduced KANs, a novel class of neural networks that have since been adopted as an alternative to MLPs. The formulation corresponding to Eq. (10) for the original implementation of KANs, known as “vanilla” KANs, is given by
@@
is a residual function and $B_{m}^{(l)}\left(\cdot\right)$ are univariate spline basis functions. The superscript $(l)$ reflects the inherent dependency of the spline basis functions on a layer-specific grid, while the subscript $m$ runs from 1 to $D=G+k$, where $G$ is the number of grid intervals and $k$ is the order of the basis functions. Comparing Eq. (13) to Eq. (10), a fundamental distinction arises between MLPs and KANs in terms of their functional representation: in MLPs, nonlinearity is introduced through fixed activation functions, while only the linear transformations between layers are trainable; in contrast, KANs replace these static activation functions with learnable ones, an aspect that highlights their potential to create more expressive architectures [KAN1, KAN3].
While vanilla KANs have demonstrated promising results in solving forward PDE problems [KINNs, Rigas], their training is burdened by the expensive computation of spline basis functions. Additionally, their dependency on a grid, although beneficial in certain applications [KAN1, KAN2, Rigas], becomes redundant when the grid remains fixed throughout training. To mitigate these inefficiencies, some approaches have introduced optimization strategies [jaxKAN, AAH3], while others have explored alternative, more computationally efficient basis functions [rbfKAN, chebyKAN, ActNet]. In this work, we adopt the latter approach and use Chebyshev-based KANs, due to their proven success in PIML [FAIR, pikanapp3, KKANs, BPIKANs]. We therefore modify the expression of Eq. (13) to
@@
where $T_{m}\left(\cdot\right)$ are Chebyshev polynomials of the first kind. The hyperbolic tangent in the argument of $T_{m}$ maps the unbounded input $x$ to the canonical range [1,1][-1,1], which is the orthogonal domain of the Chebyshev polynomials. For the purposes of the present work, the Chebyshev polynomials are explicitly defined as functions up to order $D$ to maximize computational efficiency.
@@ 3 KAN Initialization
Initialization plays a critical role in the training dynamics of deep neural networks. In the case of MLPs, extensive theoretical and empirical work has led to well-established initialization schemes tailored to different architectures and activation functions. Within the PINN framework, architectures are most often initialized using the Glorot scheme [ExpertsGuide], which aims to preserve the variance of activations and gradients across layers to avoid vanishing or exploding signals, thereby enabling stable deep learning.
For KANs – and, by extension, PIKANs – the choice of weight initialization remains largely ad hoc, with each variant in the literature adopting its own heuristic procedure. In contrast to the extensive body of work on MLP initialization, systematic studies for KANs are scarce and have thus far been limited to the vanilla formulation with B-spline bases [MOSS], where power-law and LeCun-like schemes were proposed. In the specific case of Chebyshev-based KANs, the only initialization strategy reported in the literature [chebyKAN, KKANs] involves drawing the basis coefficients $w_{jim}$ of Eq. (15) from a truncated normal distribution with zero mean and variance defined by:
$\sigma^{2}_{\text{default}}=\frac{1}{d_{\text{I}}(D+1)},$ (18)
where $D$ is the number of basis functions and $d_{\text{I}}$ is the layer’s input dimension. Building on these early efforts, we develop a Glorot-like initialization derived from a variance-preservation analysis and formulated to be agnostic to the choice of basis functions. In the present work, we assess its performance in Chebyshev-based KANs and cPIKANs, demonstrating that it offers improved training stability and accuracy over the initialization scheme of Eq. (18).
@@ 3.1 Proposed Scheme
A central principle in initialization design, which is also the core idea behind the LeCun initialization [lecun], is to preserve the variance of the signal during the forward pass. Applying this condition to Eq. (19) gives
@@
In his original work, Glorot [Glorot] further required variance preservation during the backward pass, so that gradients propagate across layers without amplification or attenuation. Applying the same reasoning to the gradient signal through Eq. (19) leads to
@@
This expression defines a basis-agnostic Glorot-like initialization rule: $\mu_{m}^{(0)}$ and $\mu_{m}^{(1)}$ capture the effect of the chosen basis functions on variance, while $d_{\text{I}}$ and $d_{\text{O}}$ play the same role as in the original Glorot scheme. The additional factor $1/D$ accounts for the contribution of the $D$ basis terms associated with each input dimension. In the special case $\mu_{m}^{(0)}=\mu_{m}^{(1)}=1$ and $D=1$, corresponding to the MLP setting, which can be interpreted as using a single basis function, Eq. (25) reduces exactly to the standard Glorot initialization. Detailed derivations of Eqs. (21) and (23) are provided in B.1.
For practical applications, it is often convenient to introduce a multiplicative gain factor, following common practice in initialization utilities provided by deep learning frameworks such as PyTorch [pytorch]. This gain term allows for empirical correction in cases where the input distribution deviates from the unit-variance assumption. Incorporating this factor, the final initialization rule becomes